How many liters of O2 (g) are needed to react completely with 56.0 L of CH4 (g) at STP to produce CO2 (g) and H2O (g)?
Given: CH4 (g) + 2O2 (g) → CO2 (g) + H2O (g)
1
Expert's answer
2020-04-24T13:34:54-0400

According to the equation 1L of CH4 reacts with 2L of O2.

Volume of O2 = V(CH4) *2 = 56*2 = 112 L